OpenStudy (anonymous):

V=(1over3)(pi)(r^2)(h) I'm trying to solve for h.

OpenStudy (anonymous):

Do you know what r is?

OpenStudy (anonymous):

no.

OpenStudy (anonymous):

So does your problem look like this? v= \[\frac{ 1 }{ 3 }(\pi)r^2h\]

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

All you have to do is divide both sides of the equation by (1/3)(pi)(r^2), so that h is all that's left on one side:\[\frac{v}{\frac{1}{3}\pi r^2} = \frac{\frac{1}{3}\pi r^2h}{\frac{1}{3}\pi r^2}\]\[\frac{v}{\frac{1}{3}\pi r^2} = h\]\[h = \frac{3v}{\pi r^2}\]